OpenStudy (anonymous):

what is the differentiation of cos x -xe^x, if x = 1.0 what is the solution"

OpenStudy (cwrw238):

derivative is - sin x -(e^x + x e^x) = -sin x - e^x ( 1 + x) plug in x = 1 to get second part

OpenStudy (cwrw238):

= -sin 1 - e(1 + 1) ( the angle = 1 radian)